# Ensemble Kalman Filtering-Aided Variational Inference for Gaussian Process State-Space Models

### 10 Dec 2023 v1

@@ IV-B1 Sparse GPSSMs
Before introducing the utilization of EnKF, we first introduce the sparse GP [30, 45], a widely utilized technique in various GP variational approximations. This method serves as a scalable approach to model the corresponding GP component within the approximate posterior distribution (see Eq. (8)), thereby guaranteeing inherent scalability in GPSSM. The main idea of the sparse GPs is to introduce a small set of inducing points $\vec{{\mathbf{z}}}\triangleq\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and $\vec{{\mathbf{u}}}\triangleq\{{\mathbf{u}}_{i}\}_{i=1}^{M}$, $M\!\ll\!T$, to serve as the surrogate of the associated GPs, where the inducing inputs, ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}},\forall i$, are placed in the same space as the latent states ${\mathbf{x}}_{t}$, while the corresponding inducing outputs ${\mathbf{u}}_{i}\!=\!f({\mathbf{z}}_{i})$ follow the same GP prior as ${\vec{\mathbf{f}}}$. With the augmentation of the inducing points, the joint distribution of the GPSSM becomes
@@ IV-B2 ELBO for sparse GPSSM
where $\mathbf{m}\!\!=\!\![{\mathbf{m}}_{1}^{\top},...,{\mathbf{m}}_{d_{x}}^{\top}]^{\top}\in\mathbb{R}^{Md_{x}}$ and $\mathbf{S}\!\!=\!\!\operatorname{diag}(\mathbf{L}_{1}\mathbf{L}_{1}^{\top},...,\mathbf{L}_{d_{x}}\mathbf{L}_{d_{x}}^{\top})\in\mathbb{R}^{Md_{x}\times Md_{x}}$ are free variational parameters. This explicit representation of the variational distribution enables scalability through the utilization of stochastic gradient-based optimization [46], as it allows for the independence of individual GP predictions given the explicit inducing points [45]. Similarly, we assume that the variational distribution over the initial state is $q({\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{0}|{\mathbf{m}}_{0},\mathbf{L}_{0}\mathbf{L}_{0}^{\top})$, where ${{\mathbf{m}}}_{0}\in\mathbb{R}^{d_{x}}$ and lower-triangular matrix $\mathbf{L}_{0}\in\mathbb{R}^{d_{x}\times d_{x}}$ are free variational parameters of $q({\mathbf{x}}_{0})$. The variational distribution of the latent states, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, obtained by integrating out ${\mathbf{f}}_{t}$ as shown in Eq. (19), will be modeled by resorting to the EnKF technique. This modeling approach will help eliminate the need for the parameterized variational distribution, a common requirement in the existing works [25, 26, 27, 28], thus overcoming the challenges associated with optimizing a large number of parameters in black-box inference networks, as discussed in Section III.
Before elucidating the methodology of employing EnKF to eliminate the parameterized assumptions associated with $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, let us first undertake the derivation of the ELBO induced by Eq. (16) and Eq. (19). After performing certain algebraic calculations, the ELBO is obtained as
@@ IV-B3 EnVI algorithm
We next proceed to outline the approximate evaluation of the ELBO in Eq. (21) using EnKF. The main idea involves employing the filtering distribution obtained from EnKF as a substitute for the corresponding variational distribution, serving as an approximation for the true posterior distribution, $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\vec{\mathbf{y}}})$.
@@
$\bar{{\mathbf{x}}}_{t}^{n}\sim p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},\mathbf{x}_{t-1}^{n}),\ \ n=1,2,...,N,$ (22)
@@
and both approximations $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ and $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ being Gaussian, determining the functional form of a common posterior transition $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, remains a challenging task. Therefore, alternative approximation methods are required to efficiently evaluate the ELBO. One alternative manner is to let $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we then end up with the same ELBO as in [26], where no filtering or smoothing is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (21) and propose our EnKF-aided variational approximation. The main result is summarized in the following proposition.
@@
The analytical evaluation of the ELBO, $\mathcal{L}$, as illustrated in Eq. (27), circumvents the explicit evaluation of the first and fourth terms in Eq. (21), and sidesteps the optimization challenge posed by the parameterized $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Consequently, EnVI substantially improves the efficiency of model learning and inference. In addition, maximizing the ELBO in Eq. (27) can be interpreted as follows: The objective is to optimize the model parameters and variational parameters such that the GPSSM can fit the observed data well at each step (indicated by the first term); simultaneously, the KL regularization terms impose constraints to prevent model overfitting (indicated by the second and third terms).
Remark 2 (Computational Complexity of EnVI).
When the dimensionality of the latent state exceeds that of the observation dimension, i.e., $d_{x}\!>\!d_{y}$, the computational complexity of the EnVI algorithm predominantly lies in the evaluation of $N$ independent particles on the GP transition during the prediction step (cf. (22)). Recall that the number of inducing points is significantly smaller than the length of the data sequence, i.e., $M\!\ll\!T$, and assume that $M\!\geq\!d_{x}$. In this context, the computational complexity of Algorithm 1 scales as $\mathcal{O}(NTd_{x}M^{2})$. However, the computation of the $N$ particles on the GP transition can be run in parallel [43], which results in the computational complexity in real-world deployments scaling as $\mathcal{O}(Td_{x}M^{2})$.
@@ IV-C More Discussions and Insights
Indeed, there are various well-established model-based filtering and smoothing techniques that can be potentially integrated into the variational inference framework for GPSSMs. Such as KF, EKF, PF, and EnKF, to name a few [1]. Among these options, EnKF, as a Monte Carlo-based method, exhibits superior capability in handling nonlinear systems when compared to KF and EKF [44]. While analytical solutions can be obtained for nonlinear systems using e.g. EKF methods, they do not align with the principle advocated by John Tukey that an approximate solution to the right problem is worth more than a precise solution to the wrong problem [48]. In comparison to PF, EnKF may have certain limitations in dealing with strongly nonlinear and non-Gaussian systems. However, sampling in PF is generally more computationally demanding, particularly for high-dimensional systems [49], whereas EnKF can be effectively employed even in extremely high-dimensional systems [50]. Moreover, the differentiability of each operation within EnKF allows the EnVI to jointly update the parameters ${{\bm{\theta}}}$ and ${\bm{\zeta}}$ through using modern automatic differentiation tools [43]. In contrast, PF involves resampling from discrete distributions [51], which pose challenges for the utilization of the reparameterization trick [31, 52], resulting in significantly higher computational complexity for gradient computations [53].
It is also vital to emphasize that the proposed EnVI algorithm falls under the NMF category (see definition in Appendix C), given the mutual coupling between the latent state trajectory and the transition function, as evident in the variational distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ in Eq. (24). This indicates that EnVI inherits the favorable characteristics of NMF algorithms [26], which have the potential to enhance learning accuracy and address the issue of underestimating state inference uncertainty, commonly encountered in MF algorithms [32].
@@ V OEnVI: Online Learning and Inference
Up to this point, it is worth noting that in contrast to the existing inference network-based variational algorithms [25, 26, 27, 28, 36, 42, 60, 58], OEnVI is a simple and straightforward extension of EnVI, which benefits from eliminating the dependence on amortized inference networks. Previous works have typically employed inference networks that take the entire sequence of observations ${\mathbf{y}}_{1:T}$ as input, necessitating a significant amount of data for offline training. While it is conceivable to limit the input length of the inference network to a shorter sequence $l$, such as ${\mathbf{y}}_{t-l:t}$, this approach still leads to prolonged training times and higher computational requirements for optimizing the inference network parameters [60]. Moreover, storing historical inputs ${\mathbf{y}}_{t-l:t}$ adds to the storage overhead, making it problematic in situations where historical data duplication and storage are not permissible. In sharp contrast, OEnVI successfully overcomes the aforementioned challenges related to the storage and optimization of inference networks. As a result, it facilitates more efficient learning and inference processes, contributing to improved overall performance. Furthermore, OEnVI offers a more comprehensive objective by simultaneously minimizing the KL divergence, accounting for data reconstruction error balance, and applying regularization to the transition function to mitigate model overfitting. This stands in contrast to the existing approaches that sorely minimize the KL divergence between filtering and prediction distributions in the context of online learning [59].
@@ VI-B System Dynamics Learning
In showcasing the superior performance of EnVI, we compare it against several prominent competing methods, namely vGPSSM [25], VCDT [27], and AD-EnKF [53]. Our implementation of VCDT incorporates an amortized inference network to address the linear growth in the number of variational parameters. The vGPSSM method adheres to the original paper’s implementation [25], while the AD-EnKF is utilized as per the default software package available online [^3] https://github.com/ymchen0/torchEnKF. The training data sequences are generated by fixing $\sigma_{\mathrm{Q}}^{2}$ at 0.05 and systematically vary $\sigma_{\mathrm{R}}^{2}$ within the range of $\{0.008,0.08,0.8\}$. As a result, we generate three sets of $T=600$ observations each for training. To ensure a fair comparison in the latent space, we adhere to Ialongo et al. [27] and keep the emission model fixed to the true generative ones for all methods, while allowing the transition to be learned. Further details, including the description of the setup for the aforementioned algorithms, are provided in Supplement E-B and the accompanying source code. The result is depicted in Table I and visualized in Fig. 3. We can find that EnVI consistently outperforms existing approaches in terms of the latent dynamic learning performance and robustness. We next conduct two ablation studies.
@@
The primary reason for this discrepancy is the increased complexity arising from additional inference network parameters, which hinder effective training. Furthermore, the inference network-based methods are prone to convergence into various unfavorable local optima, demonstrating reduced robustness. As a consequence, the learning performance of such methods often fluctuates significantly. In contrast, EnVI inherits the benefits of EnKF and avoids the need to optimize additional variational parameters from the inference network, making it more amenable to optimization and demonstrating enhanced robustness. Our experimental findings consistently demonstrate that EnVI exhibits faster convergence compared to vGPSSM and VCDT. For instance, when considering $\sigma_{\mathrm{R}}^{2}=0.008$, EnVI achieves convergence after 300 iterations, while both vGPSSM and VCDT require more iterations, as shown in Fig. 4.
@@ E-D Online Learning and Inference with OEnVI
rr9ZpSSn3Lsmir1fJZlqWlUonKsvwWwD8ymc/nXwVBeLjf7xEKhdBut9Hr9WgmkyGEkJwsy5eHG5vN5g0AKIoCAEgkEkin0wQAfN9/cXPdheu6P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 23 Dec 2023 v2

@@ IV-B1 Sparse GPSSMs
Before introducing the utilization of EnKF, we introduce the sparse GP [34], a widely utilized technique in various GP variational approximations. This method serves as a scalable approach to model the corresponding GP component within the approximate posterior distribution (see Eq. (8)), thereby guaranteeing inherent scalability in GPSSM. The main idea of the sparse GPs is to introduce a small set of inducing points $\vec{{\mathbf{z}}}\triangleq\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and $\vec{{\mathbf{u}}}\triangleq\{{\mathbf{u}}_{i}\}_{i=1}^{M}$, $M\!\ll\!T$, to serve as the surrogate of the associated GPs, where the inducing inputs, ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}},\forall i$, are placed in the same space as the latent states ${\mathbf{x}}_{t}$, while the corresponding inducing outputs ${\mathbf{u}}_{i}\!=\!f({\mathbf{z}}_{i})$ follow the same GP prior as ${\vec{\mathbf{f}}}$. With the augmentation of the inducing points, the joint distribution of the GPSSM becomes
@@ IV-B2 ELBO for sparse GPSSM
where $\mathbf{m}\!\!=\!\![{\mathbf{m}}_{1}^{\top},\ldots,{\mathbf{m}}_{d_{x}}^{\top}]^{\top}\in\mathbb{R}^{Md_{x}}$ and $\mathbf{S}\!\!=\!\!\operatorname{diag}(\mathbf{L}_{1}\mathbf{L}_{1}^{\top},\ldots,\mathbf{L}_{d_{x}}\mathbf{L}_{d_{x}}^{\top})\in\mathbb{R}^{Md_{x}\times Md_{x}}$ are free variational parameters. This explicit representation of the variational distribution enables scalability through the utilization of stochastic gradient-based optimization [47], as it allows for the independence of individual GP predictions given the explicit inducing points [46]. Similarly, we assume that the variational distribution over the initial state is $q({\mathbf{x}}_{0})=\mathcal{N}({\mathbf{x}}_{0}|{\mathbf{m}}_{0},\mathbf{L}_{0}\mathbf{L}_{0}^{\top})$, where ${{\mathbf{m}}}_{0}\in\mathbb{R}^{d_{x}}$ and lower-triangular matrix $\mathbf{L}_{0}\in\mathbb{R}^{d_{x}\times d_{x}}$ are free variational parameters of $q({\mathbf{x}}_{0})$. The variational distribution of the latent states, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, obtained by integrating out ${\mathbf{f}}_{t}$ as shown in Eq. (19), will be implicitly modeled by resorting to the EnKF technique. This modeling approach will help eliminate the need to parameterize $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, a common requirement in the existing works [28, 29, 30, 31], thus overcoming the challenges associated with optimizing a large number of variational parameters, as discussed in Section III.
Before elucidating the methodology of employing EnKF to eliminate the parameterization for the variational distribution $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, let us first undertake the derivation of the ELBO induced by Eq. (16) and Eq. (19). After performing certain algebraic calculations, the ELBO is obtained as
@@ IV-B3 EnVI algorithm
We next proceed to outline the approximate evaluation of the ELBO in Eq. (21) using EnKF. The main idea revolves around utilizing the filtering distribution acquired from EnKF as a proxy for the corresponding variational distribution, acting as an approximation for the underlying posterior distribution, $p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\vec{\mathbf{y}}})$.
@@
$\bar{{\mathbf{x}}}_{t}^{n}\sim p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},\mathbf{x}_{t-1}^{n}),\ \ n=1,2,\ldots,N,$ (22)
@@
and both approximations $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ and $q({\mathbf{x}}_{t-1}|{\vec{\mathbf{u}}})$ being Gaussian, determining the functional form of a common posterior transition $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1},{\vec{\mathbf{u}}},{\mathbf{y}}_{1:t}),\forall t$, remains a challenging task. Therefore, alternative approximation methods are required to efficiently evaluate the ELBO. One alternative manner is to simply let $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})=p({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$, we then end up with the same ELBO as in [29], where no filtering or smoothing effect is feasible. To circumvent this limitation, we examine the difference between term 1 and term 4 in Eq. (21) and propose our EnKF-aided variational approximation. The main result is summarized in the following proposition.
@@
The analytical evaluation of the ELBO, $\mathcal{L}$, as illustrated in Eq. (27), circumvents the explicit evaluation of the first and fourth terms in Eq. (21), and sidesteps the computational challenges posed by the parameterized $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}},{\mathbf{x}}_{t-1})$. Consequently, EnVI substantially can improve the efficiency of model learning and inference. In addition, maximizing the ELBO in Eq. (27) can be interpreted as follows: The objective is to optimize the model parameters and variational parameters such that the GPSSM can fit the observed data well at each step (indicated by the first term); simultaneously, the KL regularization terms impose constraints to prevent model overfitting (indicated by the second and third terms).
Remark 2.
The computational complexity of the EnVI algorithm predominantly lies in the evaluation of $N$ independent particles on the GP transition during the prediction step (see Eq. (22)). Recall that the number of inducing points is significantly smaller than the length of the data sequence, i.e., $M\!\ll\!T$, and assume that $M\!\geq\!d_{x}$. In this context, the computational complexity of Algorithm 1 scales as $\mathcal{O}(NTd_{x}M^{2})$. In practice, $N$ is often a small number, and the computation of the $N$ particles on the GP transition can be run in parallel [44], resulting in the computational complexity in real-world deployments scaling as $\mathcal{O}(Td_{x}M^{2})$. This cost matches that of existing works. Yet, the streamlined model complexity in EnVI enhances its computational robustness and accelerates convergence compared to the existing works.
@@ IV-C More Discussions and Insights
Indeed, there are various well-established model-based filtering and smoothing techniques that can be potentially integrated into the variational inference framework for GPSSMs. Such as KF, EKF, PF, and EnKF, to name a few [2]. Among these options, EnKF, as a Monte Carlo-based method, exhibits superior capability in handling nonlinear systems when compared to KF and EKF [45]. While analytical solutions can be obtained for nonlinear systems using e.g. EKF methods, they do not align with the principle advocated by John Tukey that an approximate solution to the right problem is worth more than a precise solution to the wrong problem [50]. In comparison to PF, EnKF may have certain limitations in dealing with strongly nonlinear and non-Gaussian systems. However, sampling in PF is generally more computationally demanding, particularly for high-dimensional systems [51], whereas EnKF can be effectively employed even in extremely high-dimensional systems [52]. Moreover, the inherent differentiability of each operation within EnKF allows the EnVI to jointly update the parameters ${{\bm{\theta}}}$ and ${\bm{\zeta}}$ through using modern automatic differentiation tools [44]. In contrast, PF involves resampling from discrete distributions [53], which pose challenges for the utilization of the reparameterization trick [48, 54], resulting in significantly higher computational complexity for gradient computations [55].
It is also vital to emphasize that the proposed EnVI algorithm falls under the NMF category (see definition in Appendix C), given the exploitation of the dependencies between latent states and the transition function, as evident in the variational distribution, $q({\mathbf{x}}_{t}|{\vec{\mathbf{u}}})$ in Eq. (24). This indicates that EnVI inherits the favorable characteristics of NMF algorithms [29], which have the potential to enhance learning accuracy and address the issue of underestimating state inference uncertainty, commonly encountered in MF algorithms [35].
@@ V OEnVI: Online Learning and Inference
Up to this point, it is worth noting that in contrast to the existing inference network-based variational algorithms [28, 29, 30, 31, 39, 43, 62, 60], OEnVI is a simple and straightforward extension of EnVI, which benefits from eliminating the dependence on the additional parametric variational distributions. Previous works have typically employed inference networks that take the entire sequence of observations ${\mathbf{y}}_{1:T}$ as input, necessitating a significant amount of data for offline training. While it is conceivable to limit the input length of the inference network to a shorter sequence $l$, such as ${\mathbf{y}}_{t-l:t}$, this approach still leads to prolonged training times and higher computational requirements for optimizing the inference network parameters [62]. Moreover, storing historical inputs ${\mathbf{y}}_{t-l:t}$ adds to the storage overhead, making it problematic in situations where historical data duplication and storage are not permissible. In sharp contrast, OEnVI successfully overcomes the aforementioned challenges related to the storage and optimization of inference networks. As a result, it facilitates more efficient learning and inference processes, contributing to improved overall performance. Furthermore, OEnVI offers a more comprehensive objective by simultaneously minimizing the KL divergence, accounting for data reconstruction error balance, and applying regularization to the transition function to mitigate model overfitting. This stands in contrast to the existing approaches that sorely minimize the KL divergence between filtering and prediction distributions in the context of online learning [61].
@@ VI-B System Dynamics Learning
In showcasing the superior performance of EnVI, we compare it against several prominent competing methods, namely vGPSSM [28], VCDT [30], and AD-EnKF [55]. Our implementation of VCDT incorporates an inference network to address the linear growth in the number of variational parameters. The vGPSSM method adheres to the original paper’s implementation [28], while the AD-EnKF is utilized as per the default software package available online [^3] https://github.com/ymchen0/torchEnKF. The training data sequences are generated by fixing $\sigma_{\mathrm{Q}}^{2}$ at 0.05 and systematically vary $\sigma_{\mathrm{R}}^{2}$ within the range of $\{0.008,0.08,0.8\}$. As a result, we generate three sets of $T=600$ observations each for training. To ensure a fair comparison in the latent space, we adhere to Ialongo et al. [30] and keep the emission model fixed to the true generative ones for all methods, while allowing the transition to be learned. Further details, including the description of the setup for the aforementioned algorithms, are provided in Supplement E-B and the accompanying source code. The result is depicted in Table I and visualized in Fig. 3. We can find that EnVI consistently outperforms existing approaches in terms of the latent dynamic learning performance and robustness. We next conduct two ablation studies.
@@
The primary reason for this discrepancy is the increased model and computational complexity arising from additional inference network parameters, which hinder effective training. Furthermore, the inference network-based methods are prone to convergence into various unfavorable local optima, demonstrating reduced robustness. As a consequence, the learning performance of such methods often fluctuates significantly. In contrast, EnVI inherits the benefits of EnKF and avoids the need to optimize additional variational parameters from the inference network, making it more amenable to optimization and demonstrating enhanced robustness. Our experimental findings consistently demonstrate that EnVI exhibits significantly faster convergence compared to vGPSSM and VCDT. For instance, when considering $\sigma_{\mathrm{R}}^{2}=0.008$, EnVI achieves convergence after 300 iterations, while both vGPSSM and VCDT require much more iterations, as shown in Fig. 4.